# On the Low-Temperature MCMC threshold:

### v2

@@ 1.1.3. Sparse Regression
The studied model is a standard Gaussian version of sparse regression with a very large literature in theoretical statistics, compressed sensing, and information theory (see [Wai09, HB98, VTB19, EACP11, CCL+08, RXZ21, GZ22a] and references therein). Of course also the setting is greatly motivated by the study of sparse linear models in a plethora of applied contexts such as radiology and biomedical imaging (see e.g., [LDSP08] and references therein) and genomics [CCL+08, BBHL09].
Now for our setting it is well-known (e.g., an easy adaptation of the vanilla union bound argument in [RXZ20, Theorem 4]) that as long as $\sigma^{2}=o(k),$ MAP works if $n\geq n_{\mathrm{STATS}}$ for
@@
Yet, based on the performance of standard algorithms such as LASSO [Wai09, GZ22a] and the low-degree framework [BEAH+22], polynomial-time algorithms are expected to succeed if and only if $n\geq n_{\mathrm{ALG}}$ for
@@
By applying our black-box result we can prove that in the case of sparse regression low temperature MCMC methods are in fact achieving the algorithmic threshold, in sharp contrast with the case of sparse tensor PCA we discussed above. Specifically in Theorem 4.2 we prove as long as $n\geq n_{\mathrm{MCMC}}$ for some
@@
then for any large enough $\beta>0$ the Metropolis chain with stationary measure $\pi_{\beta}$ mixes in polynomial-time.
Importantly, low temperature MCMC methods that run for polynomially many iterations cannot be in principle approximated by low-degree polynomials. For this reason, even in light of the low-degree lower bound of [BEAH+22], we focus also on whether local Markov chains can succeed when $n=o(n_{\mathrm{ALG}})$. In Theorem 4.4 we show that as long as $k=o(p^{1/3})$ whenever $n=o(n_{\mathrm{ALG}})$ for large enough $\beta>0$ a large family of low temperature local reversible MCMC methods with stationary measure $\pi_{\beta}$ (including the Metropolis chain) mix in super-polynomial time, bringing new evidence for the polynomial-time optimality of low-degree polynomials even compared to seemingly very different classes of estimators. Similar to the case of sparse tensor PCA, this result follows by leveraging the Overlap Gap Property framework from [GZ22a], alongside a lengthy and delicate modification of a second moment method inspired by a similar one in [GZ22a]. Our result yet crucially generalizes the calculation and a similar lower bound proven by [GZ22a] which was leveraging the more restrictive assumption $k=p^{o(1)}$.
In conclusion, our work identifies for sparse regression the low-temperature MCMC threshold and proves that it coincides with the algorithmic threshold, i.e., it holds up to constants
@@ 2. Main Positive result
We now start formally presenting our results. We start presenting our main results with our general positive result for local Markov chains on optimizing certain Hamiltonian of interest. Specifically, our positive results apply for the so-called Metropolis chain (also called Metropolis–Hastings process) [MRR+53, Has70, Tie94, Jer92].
@@
For correspondence to Section 1.1, $\mathcal{X}$ would be the parameter space $\Theta,$ $H$ would correspond to the logarithm of the posterior (assuming it is non-zero on the whole parameter space) and $v$ would be the MAP estimator.
We focus on running a “local” Markov chain, called the Metropolis chain [MRR+53, Has70, Tie94, Jer92], on a (simple) graph $G$ with vertex set $\mathcal{X}$. For each $x\in\mathcal{X}$ denote the neighborhood of $x$ by $\Gamma_{G}(x)=\{y\in\mathcal{X}:xy\in E\}$. Let $\Delta$ denote the maximum degree of $G$, i.e., $\Delta=\max_{x\in\mathcal{X}}|\Gamma_{G}(x)|$. The Metropolis chain associated with graph $G$ is a random walk on $G$ with transition probabilities specified by the Hamiltonian; more precisely it updates as follows in each step.
@@
It is standard that under mild conditions (namely $G$ is finite and connected), the Metropolis chain is ergodic and converges to the target distribution $\pi_{\beta}$ under any initialization, see [LP17]. In particular, if $G$ is $\Delta$-regular then the transition of the chain can be simplified as step 1 can be removed.
@@
2.2 is visualized in Fig. 1. Our first result is to show that solely under 2.2, we can directly upper bound the mixing time of the Metropolis chain for sufficiently large values of $\beta>0$. Intuitively, the Metropolis chain can find its way towards the root $v^{*}$ of the spanning tree $T.$
@@
    The Metropolis chain can find $v^{*}$ within $O(D\Delta(\log|\mathcal{X}|+\beta\mathcal{R}_{H}))$ iterations with probability at least $2/3$.
Furthermore, for any $\beta\geq\frac{2}{\delta}(\log|\mathcal{X}|+\log(\frac{40\mathcal{R}_{H}}{\delta}))$ we have:
@@
    The Metropolis chain can find $v^{*}$ within $O(\frac{\Delta\mathcal{R}_{H}}{\delta})$ iterations with probability at least $2/3$; in particular, this holds for the Randomized Greedy algorithm (i.e., $\beta=\infty$).
@@
We highlight that an interesting aspect of Theorem 2.3 is that under 2.2 we only need the existence of one “short” ascending path towards the root for the Metropolis chain to mix fast. Note that this is analogous to optimizing convex functions using gradient descent, where the gradient flows can be viewed as (continuous) descending paths from any point in the space to the minimizer. Our result shows that, in discrete settings, the mere existence of these monotonic paths suffice to find the optimizer in polynomial time.
@@ 3. Application: Sparse Gaussian additive model
We now exploit the generality of Theorem 2.3, and apply to identify the threshold for the low temperature Metropolis chain for two important statistical estimation models: the sparse Gaussian additive model, which is a generalization of sparse tensor PCA, and sparse linear regression. We start with sparse Gaussian additive model in this section.
@@ 3.1. The model
We remark for the reader’s convenience that in Section 1.1 we specialized for simplicity the results of this section to the case of sparse tensor PCA [LZ22] where for some integer $t\geq 2$, $F(v)=v^{\otimes t}$ and in particular $f(x)=x^{t}$. Yet, our positive result applies in the more general case of the sparse Gaussian additive model described above.
@@
$\pi_{\mathrm{post}}(x)\propto\exp(\lambda H(x)),\quad\forall x\in\mathcal{X}_{p,k}$
@@ 3.2. The positive result
Then, with high probability for any $\beta$ satisfying
@@
the Metropolis chain with random initialization finds $v^{*}$ within $O(\beta\lambda k^{2}p)$ iterations with probability at least $2/3$.
Furthermore, with high probability for any $\beta$ satisfying
@@
the Metropolis chain finds $v^{*}$ within $O\left(\frac{k^{t+1}}{\ell_{p,k}^{t-1}}p\right)$ iterations with probability at least $2/3$; in particular, this holds for the Randomized Greedy algorithm (i.e., $\beta=\infty$).
@@
Notice that since $t\geq 1$, one can always choose $\beta=\lambda$ and apply Theorem 3.3. In particular, our result holds for the Metropolis chain sampling from $\pi_{\mathrm{post}}.$
@@
The first part of Theorem 3.3 requires a randomized initialization for the Metropolis chain, i.e., the initial point $X_{0}$ is chosen uniformly at random from $\mathcal{X}_{p,k}$. If we further know $\lambda=\Omega(k^{t-\frac{1}{2}}\sqrt{\log p})$ (e.g., when $k=O(\sqrt{p})$), then this assumption can be removed and the result holds under any initialization.
The proof of Theorem 3.3 is given in Section 8.1.
@@ 3.3. The negative result
It is natural to wonder whether our obtained positive results is tight. In this section, we prove that in the special case of sparse tensor PCA [LZ22], the threshold for the Metropolis chain to be mixing fast is indeed captured by the above result.
@@
$Y=\frac{\lambda}{k^{t/2}}v^{\otimes t}+W$ (3.3)
for signal $v\in\{0,1\}^{p},\|v\|_{0}=k$ and noise $W\in\mathbb{R}^{p^{t}}$ with i.i.d. $N(0,1)$ entries.
@@
for any reversible Markov chain on the graph $G_{p,k}$ with stationary distribution $\pi_{\beta}$, the mixing time of the chain is $p^{\omega(1)}$.
@@ 4.1. The model
We now focus on sparse regression. Similar to Eq. 3.1, we have $\mathcal{X}_{p,k}$ for integers $0\leq k\leq p$ with the associated Johnson graph $G_{p,k}$. Let $n$ be a positive integer and let $\sigma^{2}>0$. Following the Gaussian sparse regression setting [GZ22a], let us assume that the covariate matrix $X\in\mathbb{R}^{n\times p}$ is a random matrix with i.i.d. standard Gaussian entries. Let the noise $W\in\mathbb{R}^{n}$ be distributed according to $N(0,\sigma^{2}I)$. Assume $v^{*}$ is sampled from the uniform distribution over $\mathcal{X}_{p,k}$, and let the observations be
@@ 4.2. The positive result
We work under the high signal to noise ratio assumption for sparse regression, as described in the following assumption.
@@
Assume that for a sufficiently small constant $c>0$, it holds
$\sigma^{2}\leq c\min\left\{\frac{\log p}{\log\log p},k\right\}.$
@@
Theorem 4.2 (Main algorithmic result for sparse regression).
Consider the Sparse Linear Regression model where 4.1 holds. Assume for a sufficiently large constant $C>0$,
@@
With high probability, it holds that for any $\beta=\Omega(\log(kp)/n)$, the Metropolis chain outputs $v^{*}$ after $O\left(k^{3}p\left(\log\left(\frac{pe}{k}\right)+\beta n\right)\right)$ iterations with probability at least $3/4$. Furthermore, for any $\beta>8/C$, the Metropolis chain outputs $v^{*}$ within $O(k^{2}p)$ iterations with probability at least $2/3$; in particular, this holds for the Randomized Greedy Algorithm (i.e., $\beta=\infty$).
@@
We note that for a uniform prior on $v^{*}$, the posterior given $X$ and $Y$ as described in (4.1) corresponds to the case of $\pi_{\beta}$ where $\beta=\frac{1}{2}$. Theorem 4.2 therefore implies that the Metropolis chain can sample from the posterior of sparse regression in polynomial time whenever $C>16$ or
@@
We remark that achieving such sampling results from the posterior of high dimensional estimation tasks is a non-trivial task which has seen recent progress in non-sparse settings via diffusion models techniques [MW23]. Our work implies that sampling from the posterior of sparse estimation tasks is also possible in polynomial-time via simple MCMC methods almost all the way down to the LASSO threshold [Wai09] $n_{\mathrm{ALG}}=\Theta(k\log(p/k))$.

### v3

@@ 1.1.3. Sparse Regression
The studied model is a standard Gaussian version of sparse regression with a very large literature in theoretical statistics, compressed sensing, and information theory (see [64, 31, 62, 20, 15, 55, 26] and references therein). Of course also the setting is greatly motivated by the study of sparse linear models in a plethora of applied contexts such as radiology and biomedical imaging (see e.g., [45] and references therein) and genomics [15, 9].
Now for our setting it is well-known (e.g., an easy adaptation of the vanilla union bound argument in [54, Theorem 4]) that as long as $\sigma^{2}=o(k),$ MAP works if $n\geq n_{\mathrm{STATS}}$ for
@@
Yet, based on the performance of standard algorithms such as LASSO [64, 26] and the low-degree framework [4], polynomial-time algorithms are expected to succeed if and only if $n\geq n_{\mathrm{ALG}}$ for
@@
By applying our black-box result we can prove that in the case of sparse regression low-temperature MCMC methods are in fact achieving the algorithmic threshold, in sharp contrast with the case of sparse tensor PCA we discussed above. Specifically in Theorem 4.2 we prove as long as $n\geq n_{\mathrm{MCMC}}$ for some
@@
then for any large enough $\beta>0$ the Metropolis chain on the Johnson graph with stationary measure $\pi_{\beta}$ hits $v$ in polynomial-time, extending a result from [26] which only holds for greedy dynamics and significantly lower noise levels than our result.
Importantly, low-temperature MCMC methods that run for polynomially many iterations cannot be in principle approximated by low-degree polynomials. For this reason, even in light of the low-degree lower bound of [4], we focus also on whether local Markov chains can succeed when $n=o(n_{\mathrm{ALG}})$. In Theorem 4.5 we show that as long as $k=o(p^{1/3})$ whenever $n=o(n_{\mathrm{ALG}})$ for large enough $\beta>0$ a large family of low-temperature local reversible MCMC methods on the Johnson graph with stationary measure $\pi_{\beta}$ (including the Metropolis chain) take super-polynomial time to hit $v$, bringing new evidence for the polynomial-time optimality of low-degree polynomials even compared to seemingly very different classes of estimators. Similar to the case of sparse tensor PCA, this result follows by leveraging the Overlap Gap Property framework from [26], alongside a lengthy and delicate modification of a second moment method inspired by a similar one in [26]. Our result yet crucially generalizes the calculation and a similar lower bound proven by [26] which was leveraging the more restrictive assumption $k=p^{o(1)}$.
In conclusion, our work identifies for sparse regression the threshold for the success of a large class of low-temperature MCMC methods and proves that it coincides with the algorithmic threshold, i.e., it holds up to constants
@@ 2. Main Positive result
We now start formally presenting our results. We start with presenting our main results with our general positive result for local Markov chains on optimizing certain Hamiltonians of interest. Specifically, our positive results apply for the so-called Metropolis chain (also called Metropolis–Hastings process) [49, 32, 61, 36].
@@
(The corresponding Bayesian estimation quantities) For correspondence to Section 1.1, $\mathcal{X}$ would be the parameter space $\Theta$, $H$ would correspond to the logarithm of the posterior (assuming it is non-zero on the whole parameter space) and $v^{*}$ would be the MAP estimator.
We focus on running a “local” Markov chain, called the Metropolis chain [49, 32, 61, 36], on a (simple) graph $G$ with vertex set $\mathcal{X}$. For each $x\in\mathcal{X}$ denote the neighborhood of $x$ by $\Gamma_{G}(x)=\{y\in\mathcal{X}:xy\in E\}$. Let $\Delta$ denote the maximum degree of $G$, i.e., $\Delta=\max_{x\in\mathcal{X}}|\Gamma_{G}(x)|$. The Metropolis chain associated with graph $G$ is a random walk on $G$ with transition probabilities specified by the Hamiltonian; more precisely it updates as follows in each step.
@@
It is standard that under mild conditions (namely $G$ is finite and connected), the Metropolis chain is ergodic and converges to the target distribution $\pi_{\beta}$ under any initialization, see [40]. In particular, if $G$ is $\Delta$-regular then the transition of the chain can be simplified as step 1 can be removed. We note that the uniform random proposal of $x$ in the neighborhood of $X_{t-1}$ is common in the literature (e.g. in [67]), and contrasts with an “informed” proposal such as that in [73].
@@
2.2 is visualized in Fig. 1. Our first result is to show that solely under 2.2, we can directly upper bound the time the Metropolis chain takes to hit $v^{*}$ for sufficiently large values of $\beta>0$. Intuitively, the Metropolis chain can find its way towards the root $v^{*}$ of the spanning tree $T$.
@@
    For any $T=\omega(D\Delta\left(\log|\mathcal{X}|+\beta\mathcal{R}_{H})\right)$ it holds that for large enough $n$ the Metropolis chain after $T$ iterations is at state $v^{*}$ with probability at least $2/3$;
@@
    For any $T=\omega\left(\frac{\Delta\mathcal{R}_{H}}{\delta}\right)$ and any $\beta\geq\frac{2}{\delta}\left(\log|\mathcal{X}|+\log\left(\frac{40\mathcal{R}_{H}}{\delta}\right)\right)$, it holds that for large enough $n$ the Metropolis chain after $T$ iterations is at state $v^{*}$ with probability at least $2/3$. Moreover, for the Randomized Greedy algorithm (i.e., $\beta=\infty$), once the chain hits $v^{*}$, it remains at $v^{*}$ for all future iterations after $T$ with probability 1.
@@
(“One good path suffices”) We highlight that an interesting aspect of Theorem 2.3 is that under 2.2 we only need the existence of one “short” ascending path towards the root for the Metropolis chain to hit the root fast. Note that this is analogous to optimizing convex functions using gradient descent, where the gradient flows can be viewed as (continuous) descending paths from any point in the space to the minimizer. Our result shows that, in discrete settings, the mere existence of these monotonic paths suffice to find the optimizer in polynomial time.
Remark 2.5.
(“How high is the lower bound on $\beta$?”) While our positive result holds only for “large enough” $\beta,$ we highlight that our lower bound on the applicable values of $\beta$ so that Theorem 2.3 applies is not that restrictive. For instance, as we are about to see in two applications (in sparse PCA and sparse regression), the bound is satisfied in both examples for any $\beta$ bigger than the Bayesian optimal posterior temperature of the statistical task, and it is satisfied for certain smaller values of $\beta$ than the posterior temperature as well.
@@ 3. Application: Sparse Gaussian additive model
We now exploit the generality of Theorem 2.3, and apply to identify the threshold for the low-temperature Metropolis chain for two important statistical estimation models: the sparse Gaussian additive model, which is a generalization of sparse tensor PCA, and sparse linear regression. We start with sparse Gaussian additive model in this section.
@@ 3.1. The model
We remark for the reader’s convenience that in Section 1.1 we specialized for simplicity the results of this section to the case of sparse tensor PCA [44] where for some integer $t\geq 2$, $F(v)=v^{\otimes t}$ and in particular $f(x)=x^{t}$. Yet, our positive result applies in the more general case of the sparse Gaussian additive model described above.
@@
$\pi_{\mathrm{post}}(x)\propto\exp(\lambda H(x)),\quad\forall x\in\mathcal{X}_{p,k}$ (3.2)
@@ 3.2. The positive result
Then the following statements hold:
  1. (1)
    (Moderate $\beta$): For any $\beta$ satisfying
@@
    and any $T=\omega(\beta\lambda k^{2}p)$, the Metropolis chain with random initialization after $T$ iterations is at $v^{*}$ with probability at least $2/3-o(1)$.
  2. (2)
    (Large $\beta$): For any $\beta$ satisfying
@@
    and any $T=\omega\left(\frac{k^{t+1}}{\ell_{p,k}^{t-1}}p\right)$, the Metropolis chain with random initialization after $T$ iterations is at $v^{*}$ with probability at least $2/3-o(1)$. Moreover, for the Randomized Greedy algorithm (i.e., $\beta=\infty$), once the chain hits $v^{*}$, with high probability it remains at $v^{*}$ for all future iterations.
The proof of Theorem 3.3 is given in Section 8.1.
Remark 3.4 (Hitting time dependence on $\beta$).
We note that for the case (1), that is for moderately large values of $\beta,$ our hitting time guarantee of $v^{*}$ increases with $\beta$, but in case (2), where $\beta$ exceeds some larger value than case (1), the hitting time ceases to depend on $\beta$ and it reduces to $O\left(\frac{k^{t+1}}{\ell_{p,k}^{t-1}}p\right)$ for all $\beta$ in case (2).
Remark 3.5 (Positive result for the posterior temperature).
We note that the posterior temperature, $\beta_{\mathrm{post}}=\lambda$, by assumption of Theorem 3.3 satisfies
$\beta_{\mathrm{post}}\lambda=\lambda^{2}=\Omega\left(\frac{k^{2t-1}}{\ell_{p,k}^{2t-2}}\log p\right).$
An easy calculation shows that since $k=o(p)$ and $t\geq 2$, this satisfies the moderate $\beta$ threshold for all $k$. In particular, for this regime of $\lambda,$ the Metropolis chain on the Johnson graph sampling from $\pi_{\mathrm{post}}$ hits $v^{*}$ in polynomial-time with constant probability.
@@
(Random Initialization) The first part of Theorem 3.3 requires a randomized initialization for the Metropolis chain, i.e., the initial point $X_{0}$ is chosen uniformly at random from $\mathcal{X}_{p,k}$. If we further know $\lambda=\Omega(k^{t-\frac{1}{2}}\sqrt{\log p})$ (e.g., when $k=O(\sqrt{p})$), then this assumption can be removed and the result holds under any initialization.
@@
(On applying Theorem 2.3) We highlight at this point that the proof of Theorem 3.3 proceeds by constructing the paths to verify the assumptions of our general positive result Theorem 2.3. The proof relies on a careful first moment method argument alongside the Hanson-Wright inequality to show that for “most” binary $k$-sparse vectors $v^{\prime}$ there is an ascending polynomially-long path in the Johnson graph to the planted vector $v^{*}$, in the sense that the posterior of the sparse Gaussian additive model strictly increases along the path. In particular, for zero-temperature, our result guarantees the success of the randomized greedy algorithm on the Johnson graph for maximizing the posterior. It is important that such a path construction is only possible for “most” vectors $v^{\prime}$ leading to guarantees only when the Metropolis process is initialized randomly (see Remark 3.6). It is perhaps natural to wonder how one could use our main positive result Theorem 2.3, which assumes the existence of ascending polynomially-long paths to $v^{*}$ from all states. This is a robustness feature of our main result Theorem 2.3, as we first prove that the Metropolis process will remain in this “typical” part of the state space (where polynomially-long ascending paths to $v^{*}$ exist) for an exponential amount of time with high probability and then use our main positive result on this “truncated” state space.
@@ 3.3. The negative result
It is natural to wonder whether our obtained positive results is tight. In this section, we prove that in the special case of sparse tensor PCA [44], the threshold for the Metropolis chain to hit the planted vector $v^{*}$ fast is indeed captured by the above result.
@@
$Y=\frac{\lambda}{k^{t/2}}(v^{*})^{\otimes t}+W$ (3.4)
for signal $v^{*}\in\{0,1\}^{p},\|v^{*}\|_{0}=k$ and noise $W\in\mathbb{R}^{p^{t}}$ with i.i.d. $N(0,1)$ entries.
@@
for any reversible Markov chain on the graph $G_{p,k}$ with stationary distribution $\pi_{\beta}$, the first time the chain hits $v^{*}$ is $p^{\omega(1)}$.
@@
Remark 3.9.
(Negative result for the posterior temperature) By an easy calculation, we note that the conditions of Theorem 3.8 are satisfied with $\beta_{\mathrm{post}}=\lambda$ whenever
$\lambda=\omega\left(k^{1/2}\left(\log\left(\frac{p}{k}\right)\right)^{\frac{3t-2}{2(t-2)}}\right)=\tilde{\omega}(\lambda_{\mathrm{STATS}}(t,k)),$
where the $\tilde{\omega}$ notation hides poly-logarithmic terms in $p,k.$ Therefore, Theorem 3.8 implies that any reversible Markov chain on $G_{p,k}$ whose stationary distribution is $\pi_{\mathrm{post}}$ has at least $p^{\omega(1)}$ mixing time whenever
$\tilde{\omega}(\lambda_{\mathrm{STATS}}(t,k))=\lambda=\tilde{o}(\lambda_{\mathrm{MCMC}}(t,k)).$
@@ 4.1. The model
We now focus on sparse regression. Similar to Eq. 3.1, we have $\mathcal{X}_{p,k}$ for integers $0\leq k\leq p$ with the associated Johnson graph $G_{p,k}$. Let $n$ be a positive integer and let $\sigma^{2}>0$. Following the Gaussian sparse regression setting [26], let us assume that the covariate matrix $X\in\mathbb{R}^{n\times p}$ is a random matrix with i.i.d. standard Gaussian entries. Let the noise $W\in\mathbb{R}^{n}$ be distributed according to $N(0,\sigma^{2}I)$. Assume $v^{*}$ is sampled from the uniform distribution over $\mathcal{X}_{p,k}$, and let the observations be
@@ 4.2. The positive result
We work under the high signal to noise ratio assumption for sparse regression, as described in the following (mild) assumption.
@@
Assume that
$\sigma^{2}\leq\frac{k}{\log p}.$
@@
Theorem 4.2 (Sparse Regression: Positive Result).
Consider the Sparse Linear Regression model where 4.1 holds. Assume that for a sufficiently large constant $C>0$,
@@
Then the following statements hold:
  1. (1)
    (Moderate $\beta$): For any $\beta=\Omega(\log(kp)/n)$, and any $T=\omega\left(k^{3}p\left(\log\left(\frac{pe}{k}\right)+\beta n\right)\right)$, the Metropolis chain with any initialization after $T$ iterations is at $v^{*}$ with probability at least $2/3-o(1)$.
  2. (2)
    (Large $\beta$): For any $\beta>8/C$, and any $T=\omega(k^{2}p)$, the Metropolis chain with any initialization after $T$ iterations is at $v^{*}$ with probability at least $2/3-o(1)$. Moreover, for the Randomized Greedy algorithm (i.e., $\beta=\infty$), once the chain hits $v^{*}$, with high probability it remains at $v^{*}$ for all future iterations.
@@
(Hitting time dependence on $\beta$) We note that the hitting time of the chain for moderate $\beta$ has a positive dependence on $\beta n$, but once $\beta$ enters the large $\beta$ regime, this dependence disappears, and the hitting time reduces to $O(k^{2}p)$.
Remark 4.4.
(Positive result for the posterior temperature) We note that for a uniform prior on $v^{*}$, the posterior given $X$ and $Y$ as described in (4.1) corresponds to the case of $\pi_{\beta}$ where $\beta=\beta_{\mathrm{post}}=\frac{1}{2}$. Theorem 4.2 therefore implies that the Metropolis chain can sample from the posterior of sparse regression in polynomial time whenever $C>16$ or
@@
We remark that achieving such sampling results from the posterior of high dimensional estimation tasks is a non-trivial task which has seen recent progress in non-sparse settings via diffusion models techniques [51]. Our work implies that sampling from the posterior of sparse estimation tasks is also possible in polynomial-time via simple MCMC methods almost all the way down to the LASSO threshold [64] $n_{\mathrm{ALG}}=\Theta(k\log(p/k))$.